OpenStudy (anonymous):

hi how do you write -x+3y=3 in slope intercept form i keep getting y=x+0

OpenStudy (anonymous):

Well, lets first get y onto the left side of the equation by itself. Let's add x to both sides. 3y = 3 + x Let's go ahead and switch the x and 3 to make it look more like the form y = mx + b (Slope-Intercept Form) 3y = x + 3 Almost done. Do you have any questions so far?

OpenStudy (anonymous):

do u keep the 3y together i keep just subtracting 3 from both sides

OpenStudy (anonymous):

and thats how i keep getting y=x+0

OpenStudy (anonymous):

Oh no. You cannot subtract 3 when it is being multiplied. You must apply opposite's to get rid of partial terms. Ex. 1: x + 1 = 2, to get x by itself we apply the opposite idea to the 1. Since it is +1, we must -1 from both sides. Ex. 2: 2x = 1, to get x by itself we apply the opposite of the multiplication, which is division. So we divide by two on both sides.

OpenStudy (anonymous):

So in this case... 3y = x + 3 To get y by itself - See how it is being multiplied by the 3? You must "Divide" (Opposite of the Multiplication) on both sides, yielding, x + 3 y = ------- 3

OpenStudy (anonymous):

so y=x+1?

OpenStudy (jdoe0001):

\(\bf -x+3y=3\qquad {\color{brown}{ +x}} \\ \quad \\ \cancel{ -x+x }+3y=3+x\qquad {\color{brown}{ \div 3}} \\ \quad \\ \cfrac{\cancel{ 3 }y}{\cancel{ 3 }}=\cfrac{3+x}{3}\implies y=\cfrac{3}{3}+\cfrac{x}{3}\implies y=3+\cfrac{x}{3} \\ \quad \\ y=\cfrac{x}{3}+3\implies y=\cfrac{1}{3}x+3\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw